OpenStudy (andrea14158):

Myra bought a pair of jeans that cost $55. The sales tax rate is 10 percent. What is the total amount she paid for the jeans?

OpenStudy (princeharryyy):

55 + 10%of 55 = 55+5.5 =60.5